# Ionization and temperature measurements in warm dense copper using x-ray absorption spectroscopy

### v1

@@ IV.2 Extracting Inferred Temperature and Ionization from the Cu K-edge
The measured Cu XAS from these four experimental configurations are shown in Figure 3 denoted throughout as undriven, low-drive, mid-drive, and high-drive configurations corresponding to no laser drive, 3ns drive, 2ns drive, and 1ns drive, respectively. Here, we can qualitatively see the changing plasma conditions with drive intensity. Sensitivity to temperature in the K-edge and ionization in the resonance absorption show changes in the Cu plasma conditions as the drive is changed.
@@
Analyzing the slope of the K-edge and broadening of the 1s-3p absorption features require an accurate instrument response function (IRF) to account for the spectral resolution of the measurement. The IRF was determined from the undriven configuration that measured the cold Cu K-edge. The total instrument response here includes the inherent instrument resolution as well as source broadening from the backlighter size, in the case of absorption, and from the Cu target size, in the case of K and K emission. The backlighter source function can also be determined from this measured data, since the cold Cu density and thickness are known, which are important for characterizing and analyzing the driven shots. Figure 4(a) shows the undriven measurement compared to a tabulated cold edge [31].
@@
Before the temperature and ionization can be inferred from the K-edge region, the signal contributions from the background laser drive and backlighter are removed. The wide energy range of EFX allows for the measured signal away from spectral features to be used to fit a combined background and source function consistent with our undriven measurements. The measured absorption around the K-edge is modeled by fitting to two major components: a Fermi-Dirac distribution that can capture the temperature sensitivity at the K-edge, and a set of 1s3p absorption lines from each contributing charge state to capture the absorption feature below the K-edge energy. Our fitting model utilizes a Markov Chain Monte Carlo (MCMC)[32] sampler to determine the values of the free parameters of interest to infer temperature and ionization, along with statistical uncertainties and correlations.
@@
Inferring the temperature from the Fermi-Dirac component requires four total parameters, three to position it at the K-edge (an amplitude , a centering value , and a y-offset ) and finally, a temperature to fit the K-edge slope. The absolute position of the K-edge requires knowledge of each charge state’s 1s binding energy, the chemical potential of the system, and the IPD, and therefore requires a full physics model, but the temperature can readily be determined by this simple model regardless of the absolute position of the K-edge.
@@
We fit the charge state distribution using four free parameters by assuming a mean ionization from which a super Gaussian can approximate the weighted contributions from the charge states surrounding . The MCMC model can then adjust all 30 charge state weights by adjusting the Gaussian variables (amplitude , variance , centering or , and the Gaussian exponent ). To model the 1s3p absorption feature, we used Voigt profiles for each charge state, weighted by the relative population of each state as determined by the Gaussian weights. The Voigt width is calculated with both Stark and natural broadening, which will depend on the density and ; denoted as below. The Stark broadening contribution is calculated using a model derived from Griem[33] and natural widths have been estimated from Cretin[34] calculations but were found to be negligible in comparison to the IRF and Stark broadening of the absorption feature.
@@
To calculate terms in the model such as the PPS or broadening that required an electron density, we used the density in combination with the found and this is varied within the model as a fitted solution of is converged to. There is no direct measure of density for this experimental data, however, our HYDRA simulations estimate the achieved densities of each drive, albeit with large uncertainties. The laser energy variation between our VISAR experiments to the four presented here (low–high drives and undriven) was also used to estimate the change in the simulated density values found with HYDRA. For these experiments, the laser intensities varied < 10% from the expected intensities and this variation in combination with potential timing uncertainty in the backlighter, 20 ps, is used to quantify the expected uncertainty in the simulated densities. HYDRA’s underlying EOS tables were investigated, and we found utilizing either the LEOS or Sesame EOS tables[35] does not incur a notable difference in the final densities found. The densities found via HYDRA simulations are 243, 243, and 184 g/cm3 for the low, mid, and high drives, respectively. Finally, a contribution from the K fluorescence emission that also depends on the charge state distribution is incorporated into the model. The k emission amplitude is scaled from measured from the K emission on EFX using / ratios previously verified experimentally[36]. The full fitting model then takes the form:
@@
Together we have a means to infer the temperature and ionization of the buried Cu plasma without for a forward fitting physics model. Application of this fitting method yields inferred temperatures for the low, mid, and high drives of 10.60.4, 15.70.6, and 20.80.7 eV, respectively. Figure 5(a) also shows the variation in the inferred charge state distribution by model for the mid-drive. From the resonance absorption component, we find an average ionization of 4.20.8, 5.90.8, and 6.90.8 with ad-hoc PPS for the low, mid, and high drives, respectively. In comparison, FAC, with an applied Stewart-Pyatt potential as the plasma screening model, yields similar results shown in Figure 5(a) which is compared with the FAC isolated atom calculation in black. FAC with the Stewart-Pyatt potential model estimates the strongest IPD effect on the resultant CSD for the mid- and high-drives and slightly weaker shifts in the low-drive. The found for this PPS model is 4.00.8, 6.90.8, and 8.60.8 for the low-, mid-, and high-drives, respectively. Together with our density estimate, we have an inferred set of , T, and for the XAS measurements and provide important constraints in phase space for modeling these plasmas.
In the measured spectra, we observe blue shifts of the K-edge by 7 to 30 eV (low to high drive) and shifts of 4 to 26 eV in the bound resonance 1s3p absorption, with respect to the cold edge. Figure 5(b) shows the three experimental K-edge shifts as a function of their inferred (Nguyen). The combined findings of these shifts with our inference of temperature, density, and the CSD provide a condensed representation of the experimental findings for comparisons, modeling, and further experimental work in this regime. Table 1 summarizes the inferred findings from the three drive conditions. From the measured XAS data alone we cannot discriminate which model best represents the achieved plasma conditions, but future models can forward fit the spectra and future experiments could incorporate a secondary measurement like x-ray fluorescence to constrain the ionization model further. While both approaches predict a similar CSD, within the uncertainties of the experiments, neither FAC plasma screening approach can match the edge energy, . Potentially, with more information one could separate the K-edge into component contributions from the binding energies, the chemical potential, and the IPD. To achieve this the FAC (or chosen model) calculations of the binding energies need to be experimentally validated and an accurate inference of the chemical potential is needed, most accessible through measurement of the electron density. Future work to add more experimental measurements in this regime, especially with additional constraining measurements like density, will be invaluable to improving the predictive capabilities of ionization models in the WDM regime.
@@ V Conclusions
We have presented experimental results to infer ionization and temperature by measuring high-quality XAS of uniform warm dense Cu plasmas with minimal spatial and temporal gradients. Through variation of the laser drive intensity, we were able to observe sensitivity of the K-edge region to changes in temperature and ionization. Over three experiments at varied drive conditions we inferred temperatures of 10.6 0.4, 15.7 0.6, 20.80.7 and average ionization of 4.0 0.8, 5.9 0.8, and 6.9 0.8, with Nguyen PPS, for our low-, mid-, and high-drive configurations, respectively.
The temperature and ionization conditions were inferred utilizing a functional fit to the K-edge and the bound-bound resonance absorption feature. Under these degenerate conditions, the K-edge is assumed to be dominantly shaped by the Fermi-Dirac distribution of the electron population, from which the temperature is directly inferred from the slope. This temperature inference is an essential constraint on any study of ionization in this regime. We also demonstrated fitting individual charge state contributions to the bound-bound resonance absorption feature to infer the plasma’s charge state distribution. The exact model choice leads to differences in the final inferred value but is valuable bound for comparison with more sophisticated models attempting to recreate the spectra.
Because of its simplicity, this experimental configuration is easily configurable to explore more mid-z materials or complimentary diagnostics. Future iterations can be adapted for better constraining the plasma conditions by adding additional spectroscopic diagnostics, like x-ray fluorescence spectroscopy. In conjunction with XAS the x-ray fluorescence can probe the ionic structure via emission spectra. Additionally, source size broadening was a limiting factor for the spectral resolution of these measurements and improvements could yield better insight into broadening effects. Continued experimental measurements of partially ionized metals in the warm dense regime will be essential for resolving the current discrepancies between models of the ionization potential depression and other dense effects.

### v2

@@ IV.2 Extracting Inferred Temperature and Ionization from the Cu K-edge
The measured Cu XAS from these four experimental configurations are shown in Figure 5 denoted throughout as undriven, low-drive, mid-drive, and high-drive configurations corresponding to no laser drive, 3ns drive, 2ns drive, and 1ns drive, respectively. These spectra qualitatively illustrate how the plasma conditions change, and therefore the K-edge, with drive intensity. Sensitivity to temperature in the K-edge and ionization in the resonance absorption show changes in the Cu plasma conditions as the drive is changed. We also note the presence of K fluorescence pumped by the laser driven plasma, most notably in the high-drive case. Because the source (the ablation plasma) pumps the fluorescence from a different location and early in time, it is not expected to significantly effect the XAS measurement, but it can contribute to the K-edge region shape and that must be considered in the analysis.
@@
Accurate analysis of the K-edge slope and broadening of the 1s-3p absorption features require an accurate instrument response function (IRF) to account for the spectral resolution of the measurement. The IRF was determined from the undriven configuration that measured the cold Cu K-edge. The total instrument response here includes the inherent instrument resolution as well as source broadening from the backlighter size, in the case of absorption, and from the Cu target size, in the case of K and K emission. The backlighter source function can also be determined from this measured data, since the cold Cu density and thickness is known, which are important for characterizing and analyzing the driven shots. Figure 6(a) shows the undriven measurement compared to a tabulated cold edge [34].
@@
Before the temperature and ionization can be inferred from the K-edge region, the signal contributions from the background laser drive and backlighter are removed. The wide energy range of EFX allows for the measured signal away from spectral features to be used to fit a combined background and source function consistent with our undriven measurements. The measured absorption around the K-edge is modeled by fitting to two major components: a Fermi-Dirac distribution that can capture the temperature sensitivity at the K-edge, and a set of 1s3p absorption lines from each contributing charge state to capture the absorption feature below the K-edge energy. Our fitting model utilizes a Markov Chain Monte Carlo (MCMC)[35] sampler to determine the value of the free parameters of interest to infer temperature and ionization, along with statistical uncertainties and correlations.
@@
Inferring the temperature from the Fermi-Dirac component requires four total parameters, three to position it at the K-edge (an amplitude , a centering value , and a y-offset ) and finally, a temperature to fit the K-edge slope. The absolute position of the K-edge requires knowledge of each the 1s binding energy and the density of states for each ion along with the chemical potential of the system, and the IPD, and therefore requires a full physics model. However, because of the edge’s dependence on the density of states in addition to the vacancies given by the Fermi-Dirac distribution, at higher temperatures this simple method obtains a lower-limit on the temperature, which we report here. In contrast, the HYDRA simulations tend to estimate much higher achieved temperatures; which could be a result of several mechanisms including the specific ionization model in the code.
@@
We fit the charge state distribution using four free parameters by assuming a mean ionization from which a super Gaussian can approximate the weighted contributions from the charge states surrounding . The MCMC model can then adjust all 30 charge state weights by adjusting the Gaussian variables (amplitude , variance , centering or , and the Gaussian exponent ). To model the 1s3p absorption feature, we used Voigt profiles for each charge state, weighted by the relative population of each state as determined by the Gaussian weights. The Voigt width is calculated with both Stark and natural broadening, which will depend on the density and ; denoted as below. The Stark broadening contribution is calculated using a model derived from Griem[36] and natural widths have been estimated from Cretin[37] calculations but were found to be negligible in comparison to the IRF and Stark broadening of the absorption feature; the combined effects resulted in the fitted shaped seen in Figure 6.
@@
To calculate terms in the model such as the PPS or broadening that required an electron density, we used the density in combination with the found and this is varied within the model as a fitted solution of is converged to. There is no direct measure of density for this experimental data, however, our HYDRA simulations estimate the achieved densities of each drive, albeit with large uncertainties. The laser energy variation between our VISAR experiments to the four presented here (low–high drives and undriven) was also used to estimate the change in the simulated density values found with HYDRA. The simulations used for estimating the two-sided experiments utilize the found scalings from the VISAR but with the appropriate layers and the symmetric boundary for the counter propagating shock configuration. For these experiments, the laser varied < 10% from the expected intensities and this variation in combination with potential timing uncertainty in the backlighter, 20 ps, is used to quantify the expected uncertainty in the simulated densities. HYDRA’s underlying EOS tables were investigated, and we found utilizing either the LEOS or Sesame EOS tables[38] does not incur a notable difference in the final densities found. To assess model sensitivity, we varied additional multiplicative parameters in HYDRA, including inverse bremsstrahlung (laser absorption), electron conductivity, the electron flux limiter, electron–ion coupling, radiative opacity, and artificial viscosity. Even for large deviations from nominal values (0.5–2x nominal), the resulting densities changed by less than 1% for some parameters and by no more than 8% for the most sensitive (inverse bremsstrahlung and electron conductivity), this has been included in the uncertainty of the densities. The densities found via HYDRA simulations are 243, 243, and 184 g/cm3 for the low, mid, and high drives, respectively. Finally, there is a contribution from K fluorescence emission, which also depends on the charge state distribution; this can include emission pumped by the laser-driven plasma, as observed most clearly in the high-drive experiment. The k emission amplitude is scaled from measured from the K emission on EFX using / ratios previously verified experimentally[39]. The full fitting model then takes the form:
@@
Together we have a means to infer the temperature and ionization of the buried Cu plasma without the need for a forward fitting physics model. Application of this fitting method yields inferred temperatures for the low, mid, and high drives of 10.60.4, 15.70.6, and 20.80.7 eV, respectively. Figure 7(a) also shows the variation in the inferred charge state distribution by model for the mid-drive. From the resonance absorption component, we find an average ionization of 4.20.8, 5.90.8, and 6.90.8 with ad-hoc PPS for the low, mid, and high drives, respectively. In comparison, FAC, with an applied Stewart-Pyatt potential as the plasma screening model, yields similar results shown in Figure 7(a) which is compared with the FAC isolated atom calculation in black. FAC with the Stewart-Pyatt potential model estimates the strongest IPD effect on the resultant CSD for the mid- and high-drives and slightly weaker shifts in the low-drive. The found for this PPS model is 4.00.8, 6.90.8, and 8.60.8 for the low-, mid-, and high-drives, respectively. Together with our density estimate, we have an inferred set of , T, and for the XAS measurements and provide important constraints in phase space for modeling these plasmas.
Importantly, with this modeling approach we can estimate the plasma conditions, subject to model-dependent uncertainties, and thereby place meaningful constraints on the treatment of density-dependent effects and outer-shell ionization in warm dense copper. In this regime, where the thermal, Coulomb, and Fermi energies are of comparable magnitude, a self-consistent description of bound states and the continuum remains a significant challenge for current warm dense matter models.
In the measured spectra, we observe systematic blue shifts of the K-edge by 7 to 30 eV (from low to high drive) and shifts of 4 to 26 eV in the bound 1s3p resonance relative to the cold edge. Figure 7(b) shows the three experimental K-edge shifts as a function of their inferred (Nguyen). Taken together, these measured shifts, combined with the inferred temperature, density, and charge-state distribution (CSD), provide a compact and experimentally anchored data set for direct comparison with ionization and opacity models, and for benchmarking future theoretical and experimental studies in this regime. Table 1 summarizes the inferred plasma conditions for the three drive configurations.
From the measured XAS data, we identify the plasma conditions that best represent the experiment while explicitly accounting for the influence of model-dependent effects on the inferred parameters. Although both FAC screening approaches predict similar CSDs within experimental uncertainties, neither reproduces the measured edge energy, . This discrepancy, while not unexpected, highlights specific limitations in current treatments of continuum lowering and electronic structure at these densities and temperatures. With additional experimental constraints, it may be possible to decompose the K-edge into distinct contributions from bound-state energies, the chemical potential, and ionization potential depression (IPD). Achieving this would require validated calculations of the binding energies with FAC, together with an independent and accurate determination of the electron density, which remains an experimental challenge for this platform and for WDM research more broadly.
@@ V Conclusions
We have presented experimental results inferring ionization and temperature from high-quality x-ray absorption spectroscopy (XAS) measurements of uniform warm dense copper plasmas with minimal spatial and temporal gradients. By varying the laser drive intensity, we observe clear sensitivity of the K-edge region to changes in temperature and ionization. Across three different drive conditions, we infer temperatures of 10.6 0.4 eV, 15.7 0.6 eV, 20.80.7 eV and average ionization of 4.0 0.8, 5.9 0.8, and 6.9 0.8, respectively, using the Nguyen PPS for the low-, mid-, and high-drive configurations.
The temperature and ionization of the plasma are inferred using a functional fit to the K-edge and the bound-bound resonance absorption feature. Under these degenerate conditions, the K-edge is assumed to be primarily shaped by the Fermi-Dirac distribution of the electron population, allowing the temperature to be inferred directly from the edge slope. This temperature inference provides a critical constraint for studies of ionization in this regime. We further demonstrate that fitting individual charge-state contributions to the bound-bound resonance absorption feature enables inference of the plasma charge-state distribution. While the choice of atomic model leads to variations in the inferred values, these differences provide a useful bound for comparison with more sophisticated models that currently struggle to reproduce the measured spectra. Experimental data in this WDM regime are therefore essential for improving the understanding of density-dependent effects on the atomic structure of embedded ions and for constraining future modeling efforts.
Owing to its relative simplicity, this experimental platform is readily adaptable to studies of other mid-Z materials and to the inclusion of complementary diagnostics. Future experiments could further constrain plasma conditions by incorporating additional spectroscopic measurements, such as x-ray fluorescence spectroscopy, which in conjunction with XAS can probe ionic structure through emission spectra. In the present measurements, source-size broadening limited the achievable spectral resolution, and improvements in this area may yield additional insight into line-broadening mechanisms. Continued experimental investigations of partially ionized metals in the warm dense regime will be essential for addressing persistent discrepancies in models of ionization potential depression and other dense plasma effects.
@@ VI Acknowledgements
Appendix
Appendix A Spectral Resolution of the EFX Spectrometer
Quantifying the spectral resolution of the EFX spectrometer is essential for the analysis and fitting of the copper K-edge data. Following the model outlined by Chin et al. [31] the total spectral resolution is decomposed into four contributions: source size broadening, the crystal rocking-curve contribution, detector resolution, and depth broadening due to x-ray penetration into the crystal. Each term is represented as the FWHM of a Gaussian:
Here, is the change in unit energy per distance along the detector derived from the dispersion relation. is the backlighter source size (140 m for these experiments), is the FWHM of the calculated rocking curve as a function of energy, m is the image plate resolution, and is the energy- and angle-dependent x-ray penetration depth in silicon calculated using NIST data[40]. The total spectral resolution is obtained by adding the four contributions in quadrature:
The individual contributions to the spectral resolution, as well as the total calculated resolution, are shown in Figure 8 for the EFX configured for the Cu XAS experiments at OMEGA. Source size broadening is the dominant contributor to total resolution. In the region of interest around the Cu K-edge, we find a resolution of approximately 10 eV. This is in agreement with our findings from the undriven cold Cu K-edge measurements (see Fig 6(a)).
Appendix B Experimental Timing Scan of conditions
Additional experiments were performed to verify the shock timing inferred from the one-sided VISAR measurements. In these experiments, the platform was configured in the nominal two-sided geometry (see Fig. 2). The x-ray backlighter timing was varied to measure XAS at three characteristic times: prior to shock arrival at the buried copper layer, at the nominal optimal time shortly after complete shock propagation through the copper, and at a delayed time relative to this optimum. This timing scan provides an experimental validation of the HYDRA-simulated timing used to design the primary experiments.
Figure 9 shows representative HYDRA simulations and corresponding XAS measurements from three shots with varied backlighter timing for the two-sided mid-drive configuration (2 ns laser drive). The early-time shot (blue) probes a 500 ps window before the shocks reach the buried copper layer. The nominal shot (black) is timed to the optimal window just after peak compression of the buried layer, while the late-time shot (red) is delayed by an additional 250 ps relative to the nominal timing. Figure 9(a) compares the three backlighter timing windows to the HYDRA-simulated density evolution within the buried copper layer, and Fig. 9(b) shows the corresponding measured XAS in the Cu K-edge region.
Based on the HYDRA simulations, the early-time shot is expected to probe near-ambient copper conditions, which is confirmed experimentally by the observed K-edge being identical to that of the undriven case. For both the nominal and late-time shots, the simulations predict probing only warm dense copper. The close similarity of the measured XAS spectra for these two shots indicates that the nominal timing window occurs after complete shock propagation through the copper layer, as the presence of any unshocked material would produce readily identifiable changes in the absorption spectrum. In addition, the early-time shot indicates that laser-driven preheat may produce detectable K fluorescence, but does not significantly heat the buried copper layer, as no K-edge softening associated with thermal effects is observed.
Together, these measurements demonstrate that the HYDRA simulations, once adjusted to match the one-sided VISAR data, provide sufficient accuracy for timing the x-ray backlighter to probe the optimal window just after full shock traversal of the buried copper layer.